# Vectorized Computation of Euler Characteristic Functions and Transforms

### v1

@@ Abstract
The weighted Euler characteristic transform (WECT) and Euler characteristic function (ECF) have proven to be useful tools in a variety of applications. However, current methods for computing these functions are neither optimized for speed nor do they scale to higher-dimensional settings. In this work, we present a vectorized framework for computing such topological transforms using tensor operations, which is highly optimized for GPU architectures and works in full generality across geometric simplicial complexes (or cubical complexes) of arbitrary dimension. Experimentally, the framework demonstrates significant speedups (up to $180\times$) over existing methods when computing the WECT and ECF across a variety of image datasets. Computation of these transforms is implemented in a publicly available Python package called pyECT. [^1] The pyECT package is available at https://github.com/compTAG/pyECT
@@ 1 Introduction
The field of topological data analysis (TDA) utilizes invariants from topology to study the shape of data. Among the invariants employed are homology and the original topological invariant: the Euler characteristic [euler1758elementa]. In practice, a single invariant is not typically robust enough to capture the shape of data. For this reason, filtrations (i.e., nested sequences of shapes) are often constructed from data and topological invariants are then applied. This allows for significantly more information to be retained, leading to constructions such as persistence diagrams [edel2002topological] (resulting from applying homology to filtrations) and Euler characteristic functions [Serra1982] (resulting from applying the Euler characteristic to filtrations).
@@
Topological transforms [Turner2014PHT] further extend the idea of applying topological invariants to filtrations by constructing a parameterized family of filtrations and then applying topological invariants. The starting point for topological transforms is typically a geometric simplicial or cubical complex $K$ in $\mathbb{R}^{n}$. For each direction vector $\mathbf{s}\in\mathbb{S}^{n-1}$, a height filtration of $K$ is constructed and then a topological invariant is applied to each height filtration. Applying the Euler characteristic (EC) to each height filtration results in the Euler characteristic transform (ECT) [Turner2014PHT, munch2023invitation], the focus of this paper. The ECT is a complete invariant: a geometric simplicial complex $K$ in $\mathbb{R}^{n}$ can be reconstructed solely from its ECT [curry2022many, Ghrist2018]. In other words, no information is lost when passing from the simplicial complex $K$ to its ECT.
@@
A useful variant of the ECT is the weighted Euler characteristic transform (WECT) [Jiang2020TheWE]. The WECT extends the ECT to the setting of weighted geometric complexes in $\mathbb{R}^{n}$. Much like the ECT, the WECT is a complete invariant for weighted geometric complexes [Jiang2020TheWE]. Moreover, the WECT has been employed in a broad range of applications. In the field of oncology, the WECT has been used to predict survival rates of patients with glioblastoma multiforme brain tumors [crawford2020predicting]. In plant biology, the WECT has been used to classify seed phenotypes [amezquita2022measuring]. The WECT has also proven particularly adept at image classification tasks [cisewski2023weighted, Jiang2020TheWE].
Numerous software packages already exist for computing the WECT, including Demeter [amezquita2022measuring], SINATRA [crawford2020predicting], EuCalc [lebovici2024efficientcomputationtopologicalintegral], and FastTopology [laky2024]. The packages Demeter and SINATRA processes arbitrary geometric simplicial and cubical complexes in Euclidean spaces, but are not optimized for speed. The packages FastTopology and EuCalc are optimized for speed, but only apply to axis-aligned cubical complexes (i.e., grids) in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. Two other packages, [roell2023differentiable] and [saxena2025scalable], provide vectorized implementations optimal for Graphics Processing Unit (GPU) hardware. However, these packages compute an approximation of the ECT known as the differentiable Euler characteristic transform (DECT). Furthermore, [saxena2025scalable] is an optimization for the specific case of unweighted cubical complexes in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. As outlined in [roell2023differentiable], the DECT provides several advantages for computation in neural network applications. However, our method is focused on exact computation of the WECT in full generality.
Contributions
@@
In this work, we provide a fully vectorized framework for computing WECFs and WECTs. Our implementation applies in full generality to arbitrary geometric simplicial and cubical complexes in Euclidean spaces, and supports automatic gradient computations. The vectorized operations are highly optimized for fast computation on modern GPU hardware. Moreover, our implementation significantly outperforms existing software packages, including the optimized packages EuCalc and FastTopology.
@@ 2 Background
In this section, we provide necessary foundations in computational topology and tensor computation to develop the main results. For simplicity, we restrict our attention to simplicial complexes in this paper; however, our results and code work for cubical complexes with no modifications necessary.
@@ 2.1 Lower-Star Filtrations
While the primary focus of this paper is on the WECT, our algorithm applies more generally to any finite set of lower-star filtrations. Here, we provide the necessary background on lower-star filtrations.
Definition 2.1 (Lower-Star Filtrations).
Let $K$ be a simplicial complex with vertex set $K_{0}$ and let $\operatorname{sub}(K)$ denote the set of subcomplexes of $K$. A vertex filter on $K$ is a function $f\colon K_{0}\to\mathbb{R}$}. For a vertex filter $f:K_{0}\to\mathbb{R}$, the lower-star filtration induced by $f$ is the function [^2] In this paper, we use a lower case letter to denote a filter function and the same upper case letter to denote the filtration it induces.
$F:\mathbb{R}\to\operatorname{sub}(K),$
defined as
$F(t):=\{\sigma\in K\mid\max_{v\in\sigma}f(v)\leq t\}$ (1)
where the max is taken over all vertices $v$ of the simplex $\sigma$.
The main examples of lower-star filtrations used in this paper are height filtrations. To define height filtrations, we first define geometric simplicial complexes. We adopt the definition from [edelsbrunner2010computational, Chap. 3, pp. 52], and refer the reader to this source for further details.
Definition 2.2 (Geometric Simplicial Complex).
A geometric simplicial complex is a finite collection of geometric simplices $K$ in $\mathbb{R}^{n}$ that satisfies the following two conditions:
  • If $\sigma\in K$, then all faces of $\sigma$ are in $K$.
  • If $\sigma,\tau\in K$, then $\sigma\cap\tau$ is either empty, or a face of both $\sigma$ and $\tau$.
Note that any (abstract) simplicial complex has a (non-unique) geometric realization in Euclidean space, as demonstrated, for example, in [edelsbrunner2010computational, Chap. 3, pp. 53]. Given an abstract simplicial complex $K$ with underlying space $|K|$, a function $\ell:|K|\to\mathbb{R}^{n}$ that provides a geometric realization of $K$ in $\mathbb{R}^{n}$ is called an embedding of $K$ into $\mathbb{R}^{n}$.
When working with geometric realizations of abstract simplicial complexes, we often abuse notation by avoiding explicit reference to the embedding function $\ell$. Consequently, when we refer to a vertex $v\in K_{0}$, we often mean its image $\ell(v)\in\mathbb{R}^{n}$.
Definition 2.3 (Height Functions and Filtrations).
Let $K$ be a simplicial complex embedded in $\mathbb{R}^{n}$ via the map $\ell:|K|\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{\mathbb{S}}^{n-1}$} be a direction vector. The height function in direction $\mathbf{s}$ is the map $h_{\mathbf{s}}:K_{0}\to\mathbb{R}$} defined as
$h_{\mathbf{s}}(v):=\ell(v)\cdot\mathbf{s}.$
The height filtration of $K$ in direction $\mathbf{s}$, which we denote by $H_{\mathbf{s}}$, is the lower-star filtration of $K$ induced by $h_{\mathbf{s}}$.
@@ 2.2 The Weighted Euler Characteristic Transform
Definition 2.4 (Weighted Simplicial Complexes).
A weighted simplicial complex is a pair $(K,\omega)$, where $K$ is a simplicial complex and $\omega:K\to\mathbb{R}$ is a function (referred to as the weight function).
Weighted simplicial complexes are often used to model various types of data. For example, triangular meshes are widely used to model three-dimensional surfaces, with weight functions capturing scalar fields on the surface. Another example comes from image data, where grayscale images are converted into weighted simplicial complexes via the Freudenthal triangulation [cisewski2023weighted, Jiang2020TheWE] with weights derived from the pixel intensities.
Originally introduced by Euler in 1758 for convex polyhedra [euler1758elementa], the Euler characteristic (EC) is a topological invariant that readily extends to simplicial complexes. We use an extension of the EC to weighted simplicial complexes: the weighted Euler characteristic.
Definition 2.5 (Weighted Euler Characteristic).
@@
Remark 2.1.
The Euler characteristic is a special case of the WEC. For a simplicial complex $K$, its Euler characteristic is the WEC of $(K,\mathbf{1})$, where $\mathbf{1}:K\to\mathbb{R}$ is the constant function that assigns value one to each simplex.
While the WEC is a remarkably useful invariant given its simplicity, there is limited information that a single number can capture about a simplicial complex $K$. To incorporate more information, we apply the WEC to a lower-star filtration of $K$, resulting in a weighted Euler characteristic function (WECF).
Definition 2.6 (Weighted Euler Characteristic Function).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex and let $f\colon K_{0}\to\mathbb{R}$ be a vertex filter on $K$. The weighted Euler characteristic function  [^3] Weighted Euler characteristic functions are also known as weighted Euler characteristic curves. Due to their general lack of continuity, we use the term function rather than curve. of $\mathbb{K}$ induced by $f$ is the function
$\operatorname{WECF}_{f}:\mathbb{R}\to\mathbb{R}$
defined as
$\operatorname{WECF}_{f}(t):=\chi\big(F(t),\omega|_{F(t)}\big)$ (2)
where $F$ is the lower-star filtration of $K$ induced by $f$.
Definition 2.7 (Directional WECF).
Let $\mathbb{K}=(K,\omega)$ be a weighted geometric simplicial complex in $\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The directional WECF of $\mathbb{K}$ in direction $\mathbf{s}$ is the WECF of $\mathbb{K}$ induced by the height function $h_{\mathbf{s}}$.
Remark 2.2.
In image analysis tasks (and their three-dimensional analogs), there are two common choices of vertex filters considered. The first uses the pixel intensities (i.e., grayscale values) as the vertex filter. The implementations in [laky2024, wang2022gpu] compute the result of applying the (unweighted) Euler characteristic to the corresponding lower-star filtration. We refer to this simply as the ECF of an image. The second uses the pixel intensities as weights, and the height function $h_{\mathbf{s}}$ for a given direction vector $\mathbf{s}\in\mathbb{S}^{1}$ as the vertex filter. This method is implemented in [amezquita2022measuring, tang2022topological, lebovici2024efficientcomputationtopologicalintegral].
Example 2.1.
An illustration of the directional ECF and directional WECF is presented in Figure 1. The full simplicial complex, including weighted vertices, is displayed in Figure 1(c). The filtration is along direction $(1,0)$, and the corresponding ECF and WECF are included in Figure 1(d) and Figure 1(e), respectively. The ECF and WECF are the same until time 1 when the weights of the two new vertices have weights of 0.5.
@@
Figure 1: Filtration of a simplicial complex (a-c) along direction (1,0), where the weights are specified in the green zero-simplices. The weight assigned to a simplex $\sigma$ is the maximum weight of its zero-simplex faces. The corresponding ECF (d) and WECF (e) are displayed in the direction (1,0); see Equation (2). At times 1 and 2, the ECF and WECF differ due to the WECF’s use of the weighted simplices.
Typically, there is no canonical direction vector to choose for directional WECFs. For this reason, it is common to choose all direction vectors. This results in the weighted Euler characteristic transform (WECT).
Definition 2.8 (Weighted Euler Characteristic Transform).
Let $\mathbb{K}=(K,\omega)$ be a weighted geometric simplicial complex in $\mathbb{R}^{n}$. The weighted Euler characteristic transform of $\mathbb{K}$ is the function
$\operatorname{WECT}_{\mathbb{K}}:\mathbb{S}^{n-1}\times\mathbb{R}\to\mathbb{R}$
defined as
@@
2.3 Existing Frameworks
Due to its simplicity and ease of computation, the WECT and its variants are widely used in applied settings. Definition 2.8 suggests a naïve algorithm for computing a discretization of the WECT:
  1. 1.
    Choose a finite set of directions $S=\{s_{1},\ldots,s_{m}\}\subsetneq\mathbb{S}^{n-1}$ and a finite set of height thresholds $T=\{t_{1},\ldots t_{\ell}\}\subsetneq\mathbb{R}$.
  2. 2.
    Construct the simplicial complexes $\{\sigma\mid 0pt{\sigma}{s}\leq t\}$ for each $s\in S$ and $t\in T$.
  3. 3.
    Compute the WECT of each complex.
Suppose that $n=|K|$ is the number of simplices in $K$. This naïve implementation runs in $\Theta(\ell mn)$ time, This can be improved to $\Theta(m(\ell+n\log n))$ time using a sorting-based approach. There are several publicly available implementations to compute the WECT and its variants, which are described next.
Demeter
Demeter [amezquita2022measuring] is a Python package that computes the directional ECF and ECT of axis-aligned cubical complexes. Note that it incorporates an optimization of bucket sorting the filtration values of each vertex. However, this package does not work for generic simplicial complexes, and only computes the special case of the unweighted ECT.
SINATRA
The SINATRA package [tang2022topological] is an R package that computes the directional ECF and ECT through an algorithm similar to Demeter. However, unlike Demeter, SINATRA works for arbitrary simplicial complexes. Note that it does not allow for weighted complexes and is not optimized for speed.
EuCalc
EuCalc [lebovici2024efficientcomputationtopologicalintegral] is a highly optimized package that uses a preprocessing step to compute critical points of the directional ECF. This allows for efficient computation of the exact directional ECF (as opposed to a vectorization of it). The package only accepts axis-aligned cubical complexes as input, but does generalize to weighted complexes.
FastTopology
The FastTopology framework [laky2024] is a highly optimized implementation for computing the ECF of lower-star filtrations of two- and three-dimensional axis-aligned cubical complexes. While it is designed to be highly parallelizable on CPU architectures, it does not directly support vectorized computation (e.g., on a GPU) or the WECF.
GPU ECC

### v2

@@ Abstract
The weighted Euler characteristic transform (WECT) and Euler characteristic function (ECF) have proven to be useful tools in a variety of applications. However, current methods for computing these functions are either not optimized for speed or do not scale to higher-dimensional settings. In this work, we present a vectorized framework for computing such topological descriptors using tensor operations, which is highly optimized for GPU architectures and works in full generality across simplicial and cubical complexes of arbitrary dimension. Experimentally, the framework demonstrates significant speedups over existing methods when computing the WECT and ECF across a variety of two- and three-dimensional datasets. Computation of these transforms is implemented in a publicly available Python package called pyECT.
@@ 1 Introduction
The field of topological data analysis (TDA) utilizes invariants from topology to study the shape of data. Among the invariants employed are homology and the original topological invariant: the Euler characteristic [10]. In practice, a single invariant is not typically robust enough to capture the shape of data. For this reason, filtrations (i.e., nested sequences of shapes) are often constructed from data and topological invariants are then applied. This allows for significantly more information to be retained, leading to constructions such as persistence diagrams [9] and Euler characteristic functions [20].
@@
Topological transforms [23] further extend the idea of applying topological invariants to filtrations by constructing a parameterized family of filtrations and then applying topological invariants. The starting point for topological transforms is typically a geometric simplicial or cubical complex $K$ in $\mathbb{R}^{n}$. For each direction vector $\mathbf{s}\in\mathbb{S}^{n-1}$, a directional filtration of $K$ is constructed and then a topological invariant is applied to that filtration. Applying the Euler characteristic (EC) to a single directional filtration results in the Euler characteristic function (ECF), while applying it to each directional filtration in $\mathbb{S}^{n-1}$ results in the Euler characteristic transform (ECT) [23, 15]. The ECT is a complete invariant: a geometric simplicial complex $K$ in $\mathbb{R}^{n}$ can be reconstructed solely from its ECT [6, 11]. In other words, no information is lost when passing from the simplicial complex $K$ to its ECT.
@@
Useful variants of the ECF and ECT are the weighted Euler characteristic function (WECF) and the weighted Euler characteristic transform (WECT) for weighted complexes [12]. Much like the ECT, the WECT is a complete invariant for weighted geometric complexes [12]. Moreover, the WECT has been employed in a broad range of applications, including: in oncology for predicting survival rates of patients with glioblastoma multiforme brain tumors [5], in plant biology for classifying seed phenotypes [1], and in image classification tasks [3, 12]. Due to its popularity, numerous software packages already exist for computing the WECT and its variants, including those presented in [1, 5, 13, 14, 17, 18]. While many of these packages provide useful tools, they are either not optimized for speed or are limited in the types (or dimension) of complexes they can process.
@@
In this work, we provide a fully vectorized framework for computing WECFs and WECTs. Our implementation applies in full generality to arbitrary geometric simplicial and cubical complexes in Euclidean spaces. The vectorized operations are highly optimized for fast computation on modern GPU hardware, providing speedups over existing implementations.
@@ 2 Background
In this section, we provide necessary foundations in computational topology and tensor computation to develop the main results. For simplicity, we restrict our attention to simplicial complexes in this paper; however, our results and code work for cubical complexes.
We use abstract simplicial complexes throughout this work. For a (abstract) simplicial complex $K$ and for any $i\in\{0,1,\ldots,\dim(K)\}$, we denote the set of $i$-simplices of $K$ by $K_{i}$ and treat $\sigma\in K$ as a set of vertices. Furthermore, we define $k=|K|$ and $k_{i}=|K_{i}|$.
@@ 2.1 Lower-Star Filtrations
Here, we provide the necessary background on filtrations.
Definition 1 (Filtrations and Filters).
Let $K$ be a simplicial complex and let $\operatorname{sub}(K)$ denote the set of subcomplexes of $K$. A filtration of $K$ is a function $F\colon\mathbb{R}\to\operatorname{sub}(K)$} such that for every $s\leq t\in\mathbb{R}$, we have a nesting of the respective complexes: $F(s)\subseteq F(t)$. [^1] This definition generalizes to any poset as the domain of $F$ rather than $\mathbb{R}$. However, the one-parameter setting is sufficient for our purposes.
A filter on $K$ is a function $f\colon K\to\mathbb{R}$ such that the function $F:\mathbb{R}\to\operatorname{sub}(K)$ defined by
$F(t):=\{\sigma\in K\mid f(\sigma)\leq t\}$ (1)
is a filtration.
Next, we introduce a class of filtrations that are the focus of this work.
Definition 2 (Lower-Star Filtrations).
Let $K$ be a simplicial complex with vertex set $K_{0}$. A vertex filter on $K$ is a function $f\colon K_{0}\to\mathbb{R}$}. For a vertex filter $f:K_{0}\to\mathbb{R}$, the lower-star filtration induced by $f$ is the function $F:\mathbb{R}\to\operatorname{sub}(K)$, defined as
$F(t):=\{\sigma\in K\mid\max_{v\in\sigma}f(v)\leq t\}.$ (2)
The main examples of lower-star filtrations used in this paper are directional filtrations. These are defined for any (abstract) simplicial complex $K$ with a function $\ell\colon K_{0}\to\mathbb{R}^{n}$. In practice, simplicial complexes are often naturally embedded in $\mathbb{R}^{n}$ and, in this setting, the function $\ell$ simply maps each vertex to its coordinates.
Definition 3 (Vertex Height Functions and Directional Filtrations).
Let $K$ be a simplicial complex with a map $\ell:K_{0}\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The vertex height function in direction $\mathbf{s}$ is the map $h_{\mathbf{s}}:K_{0}\to\mathbb{R}$} defined as $h_{\mathbf{s}}(v):=\ell(v)\cdot\mathbf{s}.$ The directional filtration of $K$ in direction $\mathbf{s}$, denoted as $H_{\mathbf{s}}$, is the lower-star filtration of $K$ induced by $h_{\mathbf{s}}$.
@@ 2.2 The Weighted Euler Characteristic Transform
Definition 4 (Weighted Simplicial Complexes).
A weighted simplicial complex is a pair $(K,\omega)$, where $K$ is a simplicial complex and $\omega:K\to\mathbb{R}$ is a function (that we call a weight function).
Weighted simplicial complexes are often used to model various types of data. For example, triangular meshes are widely used to model three-dimensional surfaces, with weight functions capturing scalar fields on the surface. Another example comes from image data, where grayscale images are converted into weighted simplicial complexes via the Freudenthal triangulation [3, 12] with weights derived from the pixel intensities.
Originally introduced by Euler in 1758 for convex polyhedra [10], the Euler characteristic (EC) is a topological invariant that readily extends to simplicial complexes. We use an extension of the EC to weighted simplicial complexes: the weighted Euler characteristic.
Definition 5 (Weighted Euler Characteristic).
@@
Note that the Euler characteristic is a special case of the WEC, where the weight function is equal to the constant function $\mathbf{1}:K\to\mathbb{R}$. While the WEC is a remarkably useful invariant given its simplicity, there is limited information that a single number can capture about a simplicial complex. To incorporate more information, we apply the WEC to a filtration:
Definition 6 (Weighted Euler Characteristic Function).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex and let $F\colon\mathbb{R}\to\operatorname{sub}(K)$ be a filtration of $K$. The weighted Euler characteristic function [^2] Weighted Euler characteristic functions are also known as weighted Euler characteristic curves. Due to their general lack of continuity, we use the term function rather than curve. of $\mathbb{K}$ induced by $F$ is the function $\operatorname{WECF}_{F}:\mathbb{R}\to\mathbb{R}$ defined as
$\operatorname{WECF}_{F}(t):=\chi\big(F(t),\omega|_{F(t)}\big).$ (3)
When $F$ is induced from a (vertex) filter function $f$, we write $\operatorname{WECF}_{f}:=\operatorname{WECF}_{F}$.
Definition 7 (Directional WECF).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex with a function $\ell\colon K_{0}\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The directional WECF of $\mathbb{K}$ in direction $\mathbf{s}$ is the WECF of $\mathbb{K}$ induced by the height function $h_{\mathbf{s}}$.
Remark 8.
In image analysis tasks (and their three-dimensional analogs), there are two common choices of vertex filters considered. The first (implemented in [13, 24]) uses the pixel intensities as the vertex filter. It then computes the result by applying the (unweighted) Euler characteristic to the corresponding lower-star filtration. We refer to this simply as the ECF of an image. The second, which we refer to as the directional WECF of an image, uses the pixel intensities as weights and the height function $h_{\mathbf{s}}$ for a given direction vector $\mathbf{s}\in\mathbb{S}^{1}$ as the vertex filter. This method is implemented in [1, 22, 14].
Example 9.
An illustration of the directional ECF and directional WECF is presented in Figure 1. The full simplicial complex, including weighted vertices, is displayed in Figure 1(1(c)). The filtration is along direction $(1,0)$, and the corresponding ECF and WECF are included in Figure 1(1(d)) and Figure 1(1(e)), respectively. The ECF and WECF are the same until time 1 when two new vertices are introduced with weights of 0.5.
@@
Figure 1: Filtration of a simplicial complex (1(a))–(1(c)) along direction (1,0), with vertex weights shown in green discs. The weight assigned to a simplex $\sigma$ is the maximum weight of its zero-simplex faces. The corresponding ECF (1(d)) and WECF (1(e)) are displayed in the; see Equation (3). At times 1 and 2, the ECF and WECF differ due to the WECF’s use of the weighted simplices.
Typically, there is no canonical direction vector to choose when computing directional WECFs. Hence, it is common to choose all directions resulting in the weighted Euler characteristic transform (WECT).
Definition 10 (Weighted Euler Characteristic Transform).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$. The weighted Euler characteristic transform of $\mathbb{K}$ is the function $\operatorname{WECT}_{\mathbb{K}}:\mathbb{S}^{n-1}\times\mathbb{R}\to\mathbb{R}$ defined as
@@
2.3 Existing Frameworks for Computation of the ECF and (W)ECT
Due to its simplicity and ease of computation, the WECT and its variants are widely used in applied settings (resulting in many available software packages). Demeter [1] is a Python package that computes the directional ECF and ECT of axis-aligned cubical complexes. Note that it incorporates an optimization of bucket sorting the filtration values of each vertex. However, this package does not work for generic simplicial complexes, and only computes the special case of the unweighted ECT. Next, the SINATRA package [22] is an R package that computes the directional ECF and ECT through an algorithm similar to Demeter. However, unlike Demeter, SINATRA works for arbitrary, unweighted simplicial complexes. Note that as Demeter and SINATRA are not optimized for speed, they are not included in the experimental comparison. Third, EuCalc  [^3] https://github.com/HugoPasse/Eucalc [14] is a highly optimized package that uses a preprocessing step to compute critical points of the directional ECF. This allows for efficient computation of the exact directional ECF (as opposed to a vectorization of it). The package only accepts axis-aligned cubical complexes as input, but does generalize to weighted complexes. Finally, the DECT package [^4] https://github.com/aidos-lab/dect [17] is a Python package designed to compute the differentiable ECT (DECT) of simplicial complexes. While the DECT is a slightly different variant, the package also includes an implementation for computing the ECT of simplicial or cubical complexes. Although it does not allow for weighted complexes, this package is designed for GPU architectures making it an interesting point of comparison. Furthermore, [18] provides further optimizations for computing the DECT on CUDA GPU architectures, but is consequently a far less general framework.